OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1

OpenStudy (texaschic101):

do it on your own.....

OpenStudy (anonymous):

Eh ummm @Hero

OpenStudy (texaschic101):

I do not like your response to me on the last question I helped you on. So good luck on this one, but I will not help you

OpenStudy (anonymous):

Lol @texaschic101 I was kidding tho

OpenStudy (texaschic101):

I do not like someone telling me " I suck " ...even if they are kidding. So, good bye

OpenStudy (anonymous):

You're awesome and beautiful @texaschic101

OpenStudy (anonymous):

could you help me

OpenStudy (anonymous):

You do not suck @texaschic101 I suck

OpenStudy (texaschic101):

x + 3y + 2z = -16 -->(-2) 2x - y + 2z = -15 ---------------- -2x - 6y - 4z = 32 (result of multiplying by -2) 2x - y + 2z = -15 ----------------add 0 _ 7y - 2z = 17 -7y - 2z = 17 2x - y + 2z = -15 -->(-1) 2x - 2y - z = 1 --------------- -2x + y - 2z = 15 (result of multiplying by -1) 2x - 2y - z = 1 ----------------add 0 - y - 3z = 16 -y - 3z = 16 -7y - 2z = 17 -y - 3z = 16 -->(-7) -------------- -7y - 2z = 17 7y + 21z = - 112 (result of multiplying by -7) --------------add 0 + 19z = - 95 19z = -95 z = - 5 now sub -5 in for z -7y - 2z = 17 -7y -2(-5) = 17 -7y + 10 = 17 -7y = 17 - 10 -7y = 7 y = -1 now sub known values into any original equation 2x - 2y - z = 1 2x -2(-1) - (-5) = 1 2x + 2 + 5 = 1 2x + 7 = 1 2x = 1 - 7 2x = -6 x = -3 lets check our answers 2x - y + 2z = -15 2(-3) -(-1) + 2(-5) = -15 -6 + 1 - 10 = -15 -16 + 1 = -15 -15 = -15 (correct) answer : x = -3, y = -1, z = -5 or (-3,-1,-5) Learn not to be rude and you will make more friends that are willing to help you